OpenStudy (anonymous):

Simplify the following expression: 4x - 3y + 2z - 3x - y + 3z = ? Measurement is given as length by width by height.

OpenStudy (anonymous):

x -4y + 5z

OpenStudy (anonymous):

Which table best represents the steps to solve the equation x over 8 - 5 = -1? Step Justification Step 1: x over 8 = -1 + 5 The same number + 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 4 • 8 The same number 8 can be multiplied on both sides of the equation Step 4: x = 32 Multiplying 4 and 8 on the right hand side Step Justification Step 1: x over 8 = -1 - 5 The same number - 5 can be added to both sides of the equation Step 2: x over 8 = -6 Adding -1 and -5 on the right hand side Step 3: x = -6 • 8 The same number 8 can be multiplied on both sides of the equation Step 4: x = -48 Multiplying -6 and 8 on the right hand side Step Justification Step 1: x over 8 = -1 + 5 The same number - 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 4 - 8 The same number 8 can be subtracted on both sides of the equation Step 4: x = -4 Subtracting 8 from 4 on the right hand side Step Justification Step 1: x over 8 = -1 + 5 The same number + 5 can be added to both sides of the equation Step 2: x over 8 = 4 Adding -1 and 5 on the right hand side Step 3: x = 8 ÷ 4 The same number 8 can be divided on both sides of the equation Step 4: x = 2 Subtracting 8 from 4 on the right hand side